OpenStudy (anonymous):

PLEASE PLEASE HELP I WILL GIVE A MEDIAL AND FAN YOU JUST PLEASE PLEASE HELP Match each equation with its solution for n. Match Term Definition n divided by four equals eight−ninths A) one and one−half five−eighths n equals one−fourth B) three and five−ninths two−fifths n equals three−fifths C) two and four−fifths n divided by seven equals two−fifths D) two−fifths

OpenStudy (anonymous):

SOMEONE PLEASE HELP

OpenStudy (lukebluefive):

First, write the questions out: \[\frac{n}{4} = \frac{8}{9}\] \[\frac{5}{8}*n = \frac{1}{4}\] \[\frac{2}{5}*n = \frac{3}{5}\] \[\frac{n}{7} = \frac{2}{5}\]

OpenStudy (lukebluefive):

Next, write the answer choices out: \[1\frac{1}{2} = \frac{3}{2}\] \[3\frac{5}{9} = \frac{32}{9}\] \[2\frac{4}{5} = \frac{14}{5}\] \[\frac{2}{5}\]

OpenStudy (lukebluefive):

Then, try plugging in the possible answer choices to n in the questions. Start with n = 2/5, as it's the easiest. When we get to the second one, the following happens: \[\frac{5}{8} * \frac{2}{5} = \frac{10}{40} = \frac{1}{4}\] That looks like a match! Try plugging in 3/2 to the remaining three options and see which one matches.

OpenStudy (lukebluefive):

BTW, because of the following properties: \[\frac{n}{4} = n*\frac{1}{4}\] \[\frac{n}{7} = n*\frac{1}{7}\] The first and last options can be rewritten as: \[n*\frac{1}{4} = \frac{8}{9}\] \[n*\frac{1}{7} = \frac{2}{5}\]

OpenStudy (anonymous):

Thank you @LukeBlueFive

OpenStudy (lukebluefive):

You're very welcome.